OpenStudy (unklerhaukus):

Which notation style is better; A or B? (Justify your response, or provide an alternative; C)

OpenStudy (unklerhaukus):

A vector function \(f(x)\), where \(x\) is the vector \(\begin{pmatrix}\chi_0\\\chi_1\\\chi_2\end{pmatrix}\) A:\[f(\begin{pmatrix}\chi_0\\\chi_1\\\chi_2\end{pmatrix})\] B:\[f\left(\begin{pmatrix}\chi_0\\\chi_1\\\chi_2\end{pmatrix}\right)\]

OpenStudy (kmeis002):

You could always just define your vector above as \( \mathbf{x} = \left ( \begin{array}{c c c} x_0 \\ x_1 \\x_2 \end{array} \right ) \) and then use it accordingly throughout the work: \(f(\mathbf{x})\), etc

OpenStudy (unklerhaukus):

This question is about writing the function of components explicitly.

OpenStudy (kmeis002):

I see, we could always use indices for something like: \[f(\mathbf{x}) = x_0+x_1+x_2 + ...+x_n \]

OpenStudy (unklerhaukus):

Perhaps by example isn't clear enough; A: A vector function \(f\left(\begin{pmatrix}\chi_0\\\chi_1\\\chi_2\end{pmatrix}\right)=\begin{pmatrix}\chi_0+\chi_1\\\chi_2\end{pmatrix}\) B: A vector function \(f(\begin{pmatrix}\chi_0\\\chi_1\\\chi_2\end{pmatrix})=\begin{pmatrix}\chi_0+\chi_1\\\chi_2\end{pmatrix}\)

OpenStudy (kmeis002):

I see, I prefer A quite a bit

OpenStudy (unklerhaukus):

Justify...

OpenStudy (kmeis002):

No reason other then aesthetic preference. The small than large ( just really irks me

OpenStudy (unklerhaukus):

I think the small parenthesis look bad too, but i also think the doubled parenthesis looks bad

OpenStudy (unklerhaukus):

I suppose that C1: \[f\begin{pmatrix}\chi_0\\\chi_1\\\chi_2\end{pmatrix}=\begin{pmatrix}\chi_0+\chi_1\\\chi_2\end{pmatrix}\] is another option, but it could be confused with multiplying a scalar the vector.

OpenStudy (kmeis002):

If you are working in the same space, I think you could still swing: \[f(\mathbf{x}) = \left ( \begin{array}{c} x_0 + x_1 \\ x_2 \end{array} \right ) \] But if you have different input vectors for each function, it could get annoying.

OpenStudy (unklerhaukus):

I am trying to make some think like this look better

hartnn (hartnn):

\(f\left[\begin{pmatrix}\chi_0\\\chi_1\\\chi_2\end{pmatrix}\right]\) hmm...

OpenStudy (unklerhaukus):

Thanks @hartnn! ill use that (but the the other way around)

hartnn (hartnn):

looks like a vector matrix....ohh it is a vector! :3

jaynator495 (jaynator495):

to think not to long ago you didnt know hwo to latex much... or was that another mod... xD

OpenStudy (unklerhaukus):

I didn't know of \(\TeX\) before i found OpenStudy, but that was years and years ago now

jaynator495 (jaynator495):

i meant hartnn... for whatever reason you abandoned my team... i had to write the tutorial myself instead of giving credits to all those who helped that i so badly wanted to do... but oh well now...